Kate has a summer job as a swim instructor to earn her college tuition. she gets paid $8 per hour for the first 40 hours and $ 12 per hour for each hour in the week worked over 40 hours. Kate's goal is to earn at least $ 380 per week. How many hours overtime per week will she need to achive her goal?
Please help me
Thanks

Answer by checkley77(12844)

40*8+12X=380
320+12X=380
12X=380-320
12X=60
X=60/12
X=5 HOURS PER WEEK OF OVERTIME IS REQUIRED TO MEET HER GOAL.